let a = min (x²+4x+5,x belongs to R ) and b = lim??0 1-cos 2?÷?²then find the value of a-b÷a+b

Grade:12

1 Answers

Sunil Raikwar
10 years ago
a=(x+2)2+1
so minimum value of a is 1
b=lim(x tend to 0) {1-cos2x/x2}=lim(x tend to 0) {2sin2x/x2}=2lim(x tend to 0){sinx/x}=2
(a-b)/(a+b)=(1-2)/(1+2)=-1/3

Thanks & Regards
Sunil Raikwar
askIITians faculty